OpenStudy (anonymous):

Write an equation of the line containing the given point and perpendicular to the given line. Express your answer in the form y=mx + b. (3,4); 6x + y = 9 The equation of the line is y = ?

OpenStudy (anonymous):

Not the right answer...

OpenStudy (anonymous):

first, move the x on the opposite side of the y to find the slope, -6. To get the perpendicular line you want to reciprocate and take the opposite, so -(-1/6)=1/6. y=(1/6)x+b Plug in the ordered pair. 4=1/2+b 7/2=b y=x/6+(7/2)